OpenStudy (anonymous):

Write Domain/Range: f(x)= √x-1

OpenStudy (anonymous):

is the square root on the x-1 like this: sqrt(x-1)?

OpenStudy (anonymous):

yes its over both x and -1

OpenStudy (anonymous):

\[f(x)=\sqrt{x-1}\]

OpenStudy (anonymous):

you can't take the square root of a negative, so find where x-1 starts being negative, which is at zero. So x-1 = 0 x = 1 Now test values can you put a number greater than 1 in? what about 1? what about less than 1?

OpenStudy (anonymous):

Since only values less than 1 (like -2) make the value under the square root negative, then the domain is x is greater than or equal to 1.

OpenStudy (anonymous):

so would it be (1,\[\infty \])

OpenStudy (anonymous):

no, use square brackets on the 1, because that indicates including 1, like this: [1, infiity) use the parenthesis on infinity because you can't really include it.

OpenStudy (anonymous):

gotcha.. then how do i get the range?

OpenStudy (anonymous):

The easiest way is to graph it The second easiest way is to consider it a transformation of sqrt(x) The hardest way is to find the domain of the inverse On the graph, it goes up from zero. Range is the possible y values it can ever have. With the second way, you know that the range of sqrt(x) is [0, infinity) and subtracting from x shifts it horizontally so that doesn't change it. The third way is annoying, so just go with that.

OpenStudy (anonymous):

okay, so if i dont need to graph i can just make the range [0, infin)

OpenStudy (anonymous):

Yup